# Variational Inference for Sparse Poisson Regression

### v9

@@ 3.1 Laplace prior
Suppose that $y_{i}\mathop{\sim}\limits^{\mathrm{ind}}\text{Poiss}(\lambda_{i}=\exp(\beta_{0}+X_{i}{\boldsymbol{\beta}})),\quad i=1,\ldots,n$, where $\mathbf{X}_{i}=(X_{i1},\ldots,X_{ip})$ and $\boldsymbol{\beta}=(\beta_{1},\ldots,\beta_{p-1})^{\top}$}. A sparsity-enforcing prior for $\boldsymbol{\beta}$} is the Laplace prior (Seeger, 2008). Thus, we consider the independent priors $\beta_{j}\mathop{\sim}\limits^{\mathrm{ind}}\;\text{Laplace}(0,{\eta}),\;j=1,\ldots,p$. It is well-known that this prior is identical to the hierarchical model
$\beta_{j}|\tau_{j}\mathop{\sim}\limits^{\mathrm{ind}}\;{\rm N}(0,\tau_{j}),\quad\tau_{j}\mathop{\sim}\limits^{\mathrm{iid}}\;\text{ Exp}\left(\frac{{\eta}}{2}\right),\quad j=1,\ldots,p.$
@@
$\beta_{j}|\tau_{j}\mathop{\sim}\limits^{\mathrm{ind}}$ $\;{\rm N}(0,\tau_{j}),\quad\tau_{j}|{\eta}\mathop{\sim}\limits^{\mathrm{iid}}\;\text{ Exp}\left(\frac{{\eta}}{2}\right),\quad j=1,\ldots,p,$
@@
where ${\boldsymbol{\tau}}=(\tau_{0},\tau_{1},\ldots,\tau_{p})^{\top}$.
@@
and ${\rm E}_{q}({\eta})=\frac{{p+\nu-1}}{\delta+\frac{1}{2}{\sum_{j=1}^{p-1}}{\rm E}_{q}(\tau_{j})}$, for $j=1,\ldots,p$,
@@
$\bullet$ Set proper initial values for the hyper-parameters of $q(\cdot)$ functions,
@@ 3.2 Continuous spike and slab prior
$\mathbf{X}_{i}=$ $\;(1,X_{i1},\ldots,X_{ip}),\;{\boldsymbol{\beta}=(\beta_{0},\beta_{1},\ldots,\beta_{p-1})^{\top}}$
@@
$\bullet$ Set proper initial values for the hyper-parameters of $q(\cdot)$ functions,
@@ 3.3 Bernoulli sparsity-enforcing prior
$\bullet$ Set proper initial values for the hyper-parameters of $q(\cdot)$ functions,
@@ 5 Posterior predictive mass function
Using re-paramerization $\lambda_{0}=\exp\{x_{0}{\boldsymbol{\beta}}\}$, we see that
@@
$p(y_{0}|x_{0},X,y)\approx\prod_{j=1}^{p}\sum_{\gamma_{j}=0}^{1}{\rm Ber}(\gamma_{j};P_{j})\frac{1}{y_{0}!}\int\exp\{-\exp\{x_{0}\Gamma{\boldsymbol{\beta}}\}\}\exp\{y_{0}x_{0}\Gamma{\boldsymbol{\beta}}\}{\rm N}({\boldsymbol{\beta}};{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(B)},{\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(B)})\;d{\boldsymbol{\beta}}$
@@
$p(y_{0}|x_{0},X,y)$ $\approx\prod_{j=1}^{p}\sum_{\gamma_{j}=0}^{1}{\rm Ber}(\gamma_{j};P_{j})\frac{1}{y_{0}!}\int e^{-\lambda_{0\gamma}}\lambda_{0\gamma}^{y_{0}}{\rm LN}(\lambda_{0\gamma};x_{0}\Gamma{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(B)},x_{0}\Gamma{\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(B)}\Gamma x_{0}^{\top})\;d\lambda_{0\gamma}.$ (19)
@@ 6 Simulation study
For the Laplace-VB method, we set the hyper-priors $\nu=0.0001$ and $\delta=0.01$, to obtain a mean of 0.01 and variance of 1 for the parameter $\eta$, which controls the sparsity of the coefficients. The hyper-parameters ${\rho}_{1}$ and ${\rho}_{2}$ for the CS-VB and Bernoulli-VB methods are set to 1 and $\frac{1-p_{0}}{p_{0}}$ to obtain a mean $p_{0}$ for the parameters $\pi_{j}$, where $p_{0}$ is predetermined (or estimated) value of the percentage of non-sparse coefficients. Here, we estimate $p_{0}$ using the optimal LASSO method. Also, all $a_{j}$ and $b_{j}$, hyperparameters for the Bernoulli-VB method, are set to 0.01. The hyperparameter $A$ is set to 0.01 for the Laplace-VB and CS-VB methods.
@@ 6.1 Low dimensional scenario
Figures 7 and 8 show the box plots of all aforementioned criteria for all competitive methods, for the low-dimensional scenario, with $p=10$, $n=100$, and $\mathbf{z}=(1,0,1,0,1)^{\top}$. In this scenario, all models are also compared to their MCMC versions. In implementing the MCMC methods, we utilized the R packages R2OpenBUGS and rjags within R (version 4.2.1), in conjunction with OpenBUGS (version 3.2.3) and JAGS (version 4.3.1). Each chain comprised 10,000 iterations, with a burn-in of 5,000 and a thinning interval of 10.
@@
Bernoulli-VB 0.90 (0.26) 0.95 (0.14) 0.86 (0.31) 0.95 (0.14) 0.96 (0.10)
CS-VB 0.93 (0.20) 0.93 (0.20) 0.95 (0.14) 0.92 (0.22) 0.91 (0.24)
LAPLACE-VB 0.91 (0.24) 0.93 (0.20) 0.94 (0.17) 0.93 (0.20) 0.91 (0.24)
Bernoulli-MCMC 0.94 (0.17) 0.95 (0.14) 0.94 (0.17) 0.97 (0.03) 0.96 (0.10)
CS-MCMC 0.87 (0.28) 0.88 (0.25) 0.92 (0.19) 0.88 (0.25) 0.87 (0.28)
LAPLACE-MCMC 0.90 (0.20) 0.90 (0.20) 0.94 (0.06) 0.88 (0.24) 0.90 (0.20)
@@
Bernoulli-VB 0.95 (0.14) 0.91 (0.24) 0.94 (0.17) 0.91 (0.24) 0.94 (0.17)
CS-VB 0.89 (0.27) 0.92 (0.22) 0.93 (0.20) 0.95 (0.14) 0.94 (0.17)
LAPLACE-VB 0.90 (0.26) 0.91 (0.24) 0.92 (0.22) 0.95 (0.14) 0.96 (0.10)
Bernoulli-MCMC 0.97 (0.03) 0.92 (0.22) 0.96 (0.10) 0.93 (0.20) 0.95 (0.14)
CS-MCMC 0.85 (0.30) 0.90 (0.22) 0.93 (0.14) 0.92 (0.19) 0.92 (0.19)
LAPLACE-MCMC 0.87 (0.27) 0.92 (0.15) 0.94 (0.06) 0.92 (0.15) 0.94 (0.06)
@@
Figure 8: The low-dimensional scenario simulation results: the FPR (top), the FNR (middle), and the relative computation time (bottom) for different methods.
@@
where $p(\theta_{j}|y)$ is the kernel density estimator using the MCMC sample. If $q(\theta)=p(\theta|y)$ for all values of $\theta$, we have ${\rm accuracy}(q(\theta_{j}))=100\%,$ while since both $q(\cdot)$ and $p(\cdot|y)$ are probability density functions, the minimum value of ${\rm accuracy}(q(\theta_{j}))$ is zero, which occurs when $q(\cdot)$ and $p(\cdot|y)$ have separate supports. For the Bernoulli components $q(\gamma_{j})$ and $q(Z_{j})$ for Bernoulli-VB and CS-VB methods, the accuracy is computed as follows
@@
As one can see from Table 2 and Figure 6, the accuracy of the posterior approximation for the regression coefficients is high, except for zero coefficients in the Bernoulli-VB method. Note that the variational posteriors of the regression coefficients under the Bernoulli-VB method are not directly comparable to their MCMC estimated densities. In this case, we would need to compute the accuracy of the variational posterior of ${\boldsymbol{\beta}}\Gamma$. Thus, the marginal densities are either $q(\beta_{j})$ or degenerate at zero (when $\gamma_{j}=0$). Based on the criterion in Equation (21), the accuracy of the variational posterior is reduced, while in this scenario the point mass at zero offers a notably more accurate posterior. For the simulation study where the true $\beta_{j}$ is known, if we replace $p(\beta_{j}|y)$ by $\delta(\beta_{j})p(\beta_{j}|y)$ in (21), for the Bernoulli-VB method, we obtain an average accuracy of 82.89, 75.00, 82.89, 82.89, 94.74, and 82.89, for $\beta_{1},\beta_{3},\beta_{4},\beta_{5},\beta_{7}$, and $\beta_{9}$, respectively.
Table 2 also presents the accuracy for the parameters $\boldsymbol{\gamma}$} and $\mathbf{Z}$ of the VB methods against an MCMC benchmark for the Bernoulli and CS priors. As can be seen from these values, the CS-VB has higher accuracies than the Bernoulli-VB method. The accuracy for the $\pi$ hyperparameter is moderate for both Bernoulli-VB and CS-VB methods. Regarding the variance hyper-parameters of the Bernoulli-VB and Laplace-VB methods, we can see from Table 2 that the Bernoulli-VB method has lower accuracies for the variance components, which correspond with the zero coefficients compared to the other components, while the Laplace-VB method achieves almost equal accuracies for all components $\tau_{0},\ldots,\tau_{4}$.
@@ 7 Benchmark real data analysis
To examine the prediction performance of the proposed VB methods, we have considered 5 benchmark real data sets with a count response variable. It is worth noting that we do not claim that these datasets are suitable for Poisson regression. Indeed, all these datasets exhibit overdispersion, making a Negative Binomial model a strong candidate; however, our primary aim is to benchmark the performance of the proposed sparse variational Bayes algorithms within the Poisson framework. Using these challenging, real-world datasets provides a stringent test that demonstrates our method’s robustness in accurate variable selection and its computational advantages, even under realistic conditions of model misspecification.
@@
Table 3: Test relative error means (standard deviation) for 10 random partitions of 6 benchmark real data sets.
Data set LASSO SCAD Bernoulli-VB CS-VB LAPLACE-VB
affairs 0.875 (0.120) 0.870 (0.115) 0.890 (0.111) 0.880 (0.104) 0.881 (0.104)
bike sharing 0.053 (0.005) 0.059 (0.011) 0.053 (0.006) 0.060 (0.013) 0.060 (0.013)
azcabgptca 0.552 (0.029) 0.552 (0.030) 0.569 (0.027) 0.564 (0.029) 0.563 (0.027)
azdrg112 0.860 (0.024) 0.861 (0.025) 0.876 (0.028) 0.862 (0.020) 0.866 (0.022)
azpro 0.608 (0.032) 0.607 (0.032) 0.616 (0.032) 0.611 (0.034) 0.611 (0.034)
@@ 8 Concluding remarks
The VB methods proposed in this paper provide a fast Bayesian inference about the parameters of the sparse Poisson regression model. These methods are substantially faster than the MCMC methods. The proposed VB methods are used for simultaneous estimation and variable selection in the Poisson regression model. The findings indicate that the proposed Bayesian inference achieves performance equivalent to the LASSO and SCAD penalized Poisson regression models. Beyond matching predictive accuracy, the Bayesian approach offers further advantages, including direct access to highest posterior density intervals, parameter posterior distributions, and the predictive probability mass function for the response given the data. It is observed that the proposed VB methods provide a good approximation of the marginal posterior densities. A limitation of the proposed models is their sensitivity to the assumption of the Poisson distribution of the response, which fails when there is over-dispersion and under-dispersion in the response observations. Another limitation is due to the quadratic approximation used based on the idea from Jaakkola and Jordan (2000), which results in a Gaussian approximation of $q(\boldsymbol{\beta})$, which might not provide a good approximation for the posterior. The authors address all of these issues in their proposed directions for future research. A guideline for choosing the prior is given in Appendix B, based on the results of the simulation study. The codes and data for this paper are available online on GitHub at https://github.com/mortamini/VBSparsePoisson.

### v10

@@ 3.1 Laplace prior
Suppose that $y_{i}\mathop{\sim}\limits^{\mathrm{ind}}\text{Poiss}(\lambda_{i}=\exp(\beta_{0}+X_{i}{\boldsymbol{\beta}})),\quad i=1,\ldots,n$, where $\mathbf{X}_{i}=(X_{i1},\ldots,X_{i(p-1)})$ and $\boldsymbol{\beta}=(\beta_{1},\ldots,\beta_{p-1})^{\top}$}. A sparsity-enforcing prior for $\boldsymbol{\beta}$} is the Laplace prior (Seeger, 2008). Thus, we consider the independent priors $\beta_{j}\mathop{\sim}\limits^{\mathrm{ind}}\;\text{Laplace}(0,{\eta}),\;{\color[rgb]{0,0,0}j=1,\ldots,p-1}$. It is well-known that this prior is identical to the hierarchical model
$\beta_{j}|\tau_{j}\mathop{\sim}\limits^{\mathrm{ind}}\;{\rm N}(0,\tau_{j}),\quad\tau_{j}\mathop{\sim}\limits^{\mathrm{iid}}\;\text{ Exp}\left(\frac{{\eta}}{2}\right),\quad{\color[rgb]{0,0,0}j=1,\ldots,p-1}.$
@@
$\beta_{j}|\tau_{j}\mathop{\sim}\limits^{\mathrm{ind}}$ $\;{\rm N}(0,\tau_{j}),\quad\tau_{j}|{\eta}\mathop{\sim}\limits^{\mathrm{iid}}\;\text{ Exp}\left(\frac{{\eta}}{2}\right),\quad{\color[rgb]{0,0,0}j=1,\ldots,p-1},$
@@
where ${\boldsymbol{\tau}}=(\tau_{0},\tau_{1},\ldots,{\color[rgb]{0,0,0}\tau_{p-1}})^{\top}$.
@@
${\rm E}_{q}({\eta})=\frac{{p+\nu-1}}{\delta+\frac{1}{2}{\sum_{j=1}^{p-1}}{\rm E}_{q}(\tau_{j})}$, and for $j=1,\ldots,p-1$,
@@
$\bullet$ Set proper initial values for the hyper-parameters of $q(\cdot)$ functions, using an initial sparse frequentist GLM.
@@ 3.2 Continuous spike and slab prior
$\mathbf{X}_{i}=$ $\;(1,X_{i1},\ldots,{\color[rgb]{0,0,0}X_{i(p-1)}}),\;{\boldsymbol{\beta}=(\beta_{0},\beta_{1},\ldots,\beta_{p-1})^{\top}}$
@@
$\bullet$ Set proper initial values for the hyper-parameters of $q(\cdot)$ functions, using an initial sparse frequentist GLM.
@@ 3.3 Bernoulli sparsity-enforcing prior
$\bullet$ Set proper initial values for the hyper-parameters of $q(\cdot)$ functions, using an initial sparse frequentist GLM.
@@ 5 Posterior predictive mass function
Using re-parametrization $\lambda_{0}=\exp\{x_{0}{\boldsymbol{\beta}}\}$, we see that
@@
$p(y_{0}|x_{0},X,y)\approx\prod_{j=1}^{p-1}\sum_{\gamma_{j}=0}^{1}{\rm Ber}(\gamma_{j};P_{j})\frac{1}{y_{0}!}\int\exp\{-\exp\{x_{0}\Gamma{\boldsymbol{\beta}}\}\}\exp\{y_{0}x_{0}\Gamma{\boldsymbol{\beta}}\}{\rm N}({\boldsymbol{\beta}};{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(B)},{\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(B)})\;d{\boldsymbol{\beta}}$
@@
$p(y_{0}|x_{0},X,y)$ $\approx\prod_{j=1}^{p-1}\sum_{\gamma_{j}=0}^{1}{\rm Ber}(\gamma_{j};P_{j})\frac{1}{y_{0}!}\int e^{-\lambda_{0\gamma}}\lambda_{0\gamma}^{y_{0}}{\rm LN}(\lambda_{0\gamma};x_{0}\Gamma{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(B)},x_{0}\Gamma{\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(B)}\Gamma x_{0}^{\top})\;d\lambda_{0\gamma}.$ (19)
@@ 6 Simulation study
For the Laplace-VB method, we set the hyper-priors $\nu=0.0001$ and $\delta=0.01$, to obtain a mean of 0.01 and variance of 1 for the parameter $\eta$, which controls the sparsity of the coefficients. The hyper-parameters ${\rho}_{1}$ and ${\rho}_{2}$ for the CS-VB and Bernoulli-VB methods are both set to 1. Also, all $a_{j}$ and $b_{j}$, hyperparameters for the Bernoulli-VB method, are set to 0.01. The hyperparameter $A$ is set to 0.01 for the Laplace-VB and CS-VB methods.
@@ 6.1 Low dimensional scenario
Figures 7 and 8 show the box plots of all aforementioned criteria for all competitive methods, for the low-dimensional scenario, with $p=10$, $n=100$, and $\mathbf{z}=(1,0,1,0,1)^{\top}$. In this scenario, all models are also compared to their MCMC versions. In implementing the MCMC methods, we utilized the R packages R2OpenBUGS and rjags within R (version 4.2.1), in conjunction with OpenBUGS (version 3.2.3) and JAGS (version 4.3.1). Each chain comprised 10,000 iterations, with a burn-in of 5,000 and a thinning interval of 10. The Gelman-Rubin diagnostics are computed for each parameter and reported in the supplementary material. These values confirm the convergence of the MCMC algorithm for most parameters.
@@
Bernoulli-VB 0.92 (0.27) 0.94 (0.24) 0.94 (0.24) 0.96 (0.20) 0.98 (0.14)
CS-VB 0.92 (0.27) 0.96 (0.20) 0.95 (0.22) 0.96 (0.20) 0.96 (0.20)
LAPLACE-VB 0.92 (0.27) 0.95 (0.22) 0.94 (0.24) 0.96 (0.20) 0.97 (0.17)
Bernoulli-MCMC 0.96 (0.20) 0.99 (0.10) 0.98 (0.14) 0.96 (0.20) 0.99 (0.10)
CS-MCMC 0.94 (0.24) 0.97 (0.17) 0.94 (0.24) 0.97 (0.17) 0.99 (0.10)
LAPLACE-MCMC 0.95 (0.22) 0.96 (0.20) 0.96 (0.20) 0.97 (0.17) 0.97 (0.17)
@@
Bernoulli-VB 0.97 (0.17) 0.87 (0.34) 0.98 (0.14) 0.88 (0.32) 0.94 (0.24)
CS-VB 0.94 (0.24) 0.96 (0.20) 0.98 (0.14) 0.94 (0.24) 0.91 (0.29)
LAPLACE-VB 0.94 (0.24) 0.95 (0.22) 0.98 (0.14) 0.94 (0.24) 0.92 (0.27)
Bernoulli-MCMC 0.98 (0.14) 0.96 (0.20) 0.98 (0.14) 0.94 (0.24) 0.96 (0.20)
CS-MCMC 0.95 (0.22) 0.96 (0.20) 0.98 (0.14) 0.96 (0.20) 0.94 (0.24)
LAPLACE-MCMC 0.96 (0.20) 0.96 (0.20) 0.99 (0.10) 0.96 (0.20) 0.93 (0.26)
@@
Figure 8: The low dimensional scenario simulation results: the FPR (top), the FNR (middle), and the relative computation time (bottom) for different methods.
@@
where $p(\theta_{j}|y)$ is the kernel density estimator using the MCMC sample. If $q(\theta)=p(\theta|y)$ for all values of $\theta$, we have ${\rm accuracy}(q(\theta_{j}))=100\%,$ while since both $q(\cdot)$ and $p(\cdot|y)$ are probability density functions, the minimum value of ${\rm accuracy}(q(\theta_{j}))$ is zero, which occurs when $q(\cdot)$ and $p(\cdot|y)$ have separate supports.
Note that the variational posteriors of the regression coefficients under the Bernoulli-VB method are not directly comparable to their MCMC estimated densities. In this case, we would need to compute the accuracy of the variational posterior of ${\boldsymbol{\beta}}\Gamma$. Thus, the marginal densities are either $q(\beta_{j})$ or degenerate at zero (when $\gamma_{j}=0$). Based on the criterion in Equation (21), the accuracy of the variational posterior is reduced, while in this scenario the point mass at zero offers a notably more accurate posterior. Thus, for the simulation study where the true $\beta_{j}$ is known, we have replace $p(\beta_{j}|y)$ by $\delta(\beta_{j})p(\beta_{j}|y)$ in (21), for the Bernoulli-VB method.
For the Bernoulli components $q(\gamma_{j})$ and $q(Z_{j})$ for Bernoulli-VB and CS-VB methods, the accuracy is computed as follows
@@
As one can see from Table 2 and Figure 6, the accuracy of the posterior approximation for the regression coefficients is high. Table 2 also presents the accuracy for the parameters $\boldsymbol{\gamma}$} and $\mathbf{Z}$ of the VB methods against an MCMC benchmark for the Bernoulli and CS priors. As can be seen from these values, the CS-VB has higher accuracies than the Bernoulli-VB method. The accuracy for the $\pi$ hyperparameter is moderate for both Bernoulli-VB and CS-VB methods. Regarding the variance hyper-parameters of the Bernoulli-VB and Laplace-VB methods, we can see from Table 2 that the Bernoulli-VB method has lower accuracies for the variance components, which correspond with the zero coefficients compared to the other components, while the Laplace-VB method achieves almost equal accuracies for all components $\tau_{0},\ldots,\tau_{4}$.
@@ 7 Benchmark real data analysis
To examine the prediction performance of the proposed VB methods, we have considered 5 benchmark real data sets with a count response variable. It is worth noting that we do not claim that these datasets are suitable for Poisson regression. Indeed, all these datasets exhibit overdispersion, making a Negative Binomial model a strong candidate; however, our primary aim is to benchmark the performance of the proposed sparse variational Bayes algorithms within the Poisson framework. Using these challenging, real-world datasets provides a stringent test that demonstrates our method’s robustness in accurate variable selection and its computational advantages, even under realistic conditions of model misspecification. Furthermore, we have also fitted the LASSO negative binomial regression from the mpath package in R, and the results show that the Poisson regression models generally perform better than it in terms of test relative error.
@@
Table 3: Test relative error means and (standard deviations) for 10 random partitions of 6 benchmark real data sets.
Data set LASSO-NB LASSO-Poiss SCAD-Poiss Bernoulli-VB CS-VB LAPLACE-VB
affairs 0.928 0.909 0.909 0.915 0.917 0.918
(0.084) (0.086) (0.087) (0.070) (0.068) (0.065)
bike sharing 0.112 0.053 0.054 0.053 0.054 0.054
(0.024) (0.005) (0.005) (0.004) (0.005) (0.005)
azcabgptca 0.539 0.537 0.538 0.556 0.556 0.549
(0.039) (0.038) (0.038) (0.034) (0.040) (0.039)
azdrg112 0.851 0.851 0.851 0.877 0.850 0.850
(0.028) (0.028) (0.029) (0.022) (0.027) (0.027)
azprocedure 0.629 0.626 0.626 0.632 0.627 0.627
(0.026) (0.025) (0.026) (0.023) (0.025) (0.025)
azpro 0.622 0.619 0.619 0.625 0.619 0.620
(0.048) (0.048) (0.048) (0.045) (0.047) (0.047)
@@ 8 Concluding remarks
The VB methods proposed in this paper provide a fast Bayesian inference about the parameters of the sparse Poisson regression model. These methods are substantially faster than the MCMC methods. The proposed VB methods are used for simultaneous estimation and variable selection in the Poisson regression model. The findings indicate that the proposed Bayesian inference achieves performance equivalent to the LASSO and SCAD penalized Poisson regression models. Beyond matching predictive accuracy, the Bayesian approach offers further advantages, including direct access to highest posterior density intervals, parameter posterior distributions, and the predictive probability mass function for the response given the data. It is observed that the proposed VB methods provide a good approximation of the marginal posterior densities. A limitation of the proposed models is their sensitivity to the assumption of the Poisson distribution of the response, which fails when there is over-dispersion and under-dispersion in the response observations. Another limitation is due to the quadratic approximation used based on the idea from Jaakkola and Jordan (2000), which results in a Gaussian approximation of $q(\boldsymbol{\beta})$, which might not provide a good approximation for the posterior. The authors address all of these issues, including development of the VB model for the sparse negative binomial regression, in their proposed directions for future research. A guideline for choosing the prior is given in Appendix B, based on the results of the simulation study. The codes and data for this paper are available online on GitHub at https://github.com/mortamini/VBSparsePoisson.